OpenStudy (anonymous):

List all the asymptotes for the following. a) y=2/(x^2-1) y=2/((x+1)(x-1)) Vertical Asymptote: X = 1, X = -1 Horizontal Asymptote: None Oblique Asymptote: Missing Points:

OpenStudy (anonymous):

is there horizontal asymptote

OpenStudy (anonymous):

\[y = 2/(x ^{2}+1)\]

OpenStudy (anonymous):

I would say yes... :) Since the degree of denominator > nominator, you could ask yourself what happens if the limit of x->inf (+-). The function here would go to zero, and that would be the horzontal asymptote.. Hope this helps! //Ali

OpenStudy (anonymous):

No oblique one.. You find the oblique one if the Nominator is 1 degree lager than the denominator.

OpenStudy (anonymous):

also isnt degree of numerator greater since x^2 -1 = 0

OpenStudy (anonymous):

With the degre you mean the degree of the variable asked for. So no the degree of the denominator here, is greater than the nominator. Has nothing to do with x^2-1=0.